OpenStudy (anonymous):

Word Problem Need Help!!! Chen, Juanati, and Lou went to the gym together to work out. Find how much each of the kids can leg-press given the following clues. - Chen said that Juanita and Lou averaged 87 pounds . - Juanita said that Chen ;e.g.-pressed 6 pounds more than Lou. - Lou said that eight times Juanita's amount equal seven times Chen's amount.

OpenStudy (anonymous):

C = L + 6 -->L = C - 6 7C = 8J -->7/8C = J (J + L)/2 = 87 (7/8C + C - 6)/2 = 87 (7/8C + 8/8C - 6)/2 = 87 (multiply by 2) 7/8C + 8/8C - 6 = 174 15/8C = 174 + 6 15/8C = 180 C = 180 * 8/15 C = 1440/15 C = 96 C = L + 6 96 = L + 6 96 - 6 = L 90 = L 7C = 8I 7(96) = 8J 672 = 8J 672/8 = J 84 = J check.... C = L + 6 : 96 = 90 + 6 : 96 = 96 (correct) 7C = 8J : 7(96) = 8(84): 672 = 672 (correct) (J + L)/2 = 87 : (84 + 90)/2 = 87 : 174/2 = 87 : 87 = 87 (correct) Juanita can leg press 84 lbs Chen can leg press 96 lbs Lou can leg press 90 lbs